OpenStudy (anonymous):

Hanna makes her own furniture polish using 25% vinegar and 75% olive oil. Write and solve an equation to determine how much total solution she made if she used 3 cups of olive oil. Show your work.

OpenStudy (anonymous):

according to question : 75% olive oil is used and quantity is 3 cup so 75 % of x ( let x be the total Polish mixture ) =3 (75/100) * x=3 x= 3*(100/75) x=3(4/3) -----------{25 common } so x = 4 total quantity is 4 cup so venagar used will be 1 cup and olive oil 3 cup hope it helped you @JulieeeXD